OpenStudy (anonymous):

y = 5 Sin^2(x) 0 <= x <= pi ... about x-axis Find volume by rotating around the region Medal for showing me how to do this! :)

OpenStudy (anonymous):

I thought this was be appropriate for the disc method. I squared the function and multiplied by pi to solve the area. However when I get to sin^4x I get stuck. I can only think of some identities. sin(2x)..? Someone please help! :(

OpenStudy (abmon98):

\[\int\limits_{0}^{\Pi}\left(( 5\sin^2x)^2 \right)*\]

OpenStudy (abmon98):

pi

OpenStudy (anonymous):

?.. this is what I typed, I am looking to integrate and solve, if you cannot help, please don't respond :)

OpenStudy (abmon98):

5sin^2x*5sin^2x=25sin^4x^2 25sin^4(pi)^2-25sin^4(0)^2*pi

OpenStudy (anonymous):

no...